On December 31, 2017, Dow Steel Corporation had 790,000 shares of common stock and 49,000 shares of 7%, noncumulative, nonconvertible preferred stock issued and outstanding. Dow issued a 4% common stock dividend on May 15 and paid cash dividends of $590,000 and $88,000 to common and preferred shareholders, respectively, on December 15, 2018.

On February 28, 2018, Dow sold 72,000 common shares. Also, as a part of a 2017 agreement for the acquisition of Merrill Cable Company, another 19,000 shares (already adjusted for the stock dividend) are to be issued to former Merrill shareholders on December 31, 2019, if Merrill's 2019 net income is at least $690,000. In 2018, Merrill's net income was $820,000.

In keeping with its long-term share repurchase plan, 6,000 shares were retired on July 1. Dow's net income for the year ended December 31, 2018, was $3,050,000. The income tax rate is 40%.

As part of an incentive compensation plan, Dow granted incentive stock options to division managers at December 31 of the current and each of the previous two years. Each option permits its holder to buy one share of common stock at an exercise price equal to market value at the date of grant and can be exercised one year from that date. Information concerning the number of options granted and common share prices follows:

Date Granted Options Granted Share Price

(adjusted for the stock dividend)

December 31, 2016 26,000 $ 44

December 31, 2017 21,000 $ 53

December 31, 2018 24,500 $ 52

The market price of the common stock averaged $52 per share during 2018.

On July 12, 2016, Dow issued $500,000 of convertible 10% bonds at face value. Each $1,000 bond is convertible into 40 common shares (adjusted for the stock dividend).

Required: Compute Dow's basic and diluted earnings per share for the year ended December 31, 2018.
Answer #1

IF THIS ANSWER IS HELPFUL, SUPPORT ME WITH A THUMPS-UP

BASIC EARNING PER SHARE

Numerator basic EPS calculations are as:-

Net income = 3,050,000

Preferential dividend = 88,000 (because the preferred stock is non-cumulative, only include the dividend that are actually made.)

Denominator basic EPS calculations are as:-

Weighted average shares of common stock outstanding.

1/1-12/31

790,000*(12/12)=

790,000*1.04=

821600

2/28-12/31

72,000*(10/12)=

60000*1.04=

62400

7/1-12/31

(6,000)*(6/12)=

(3000)

weighted average shares

881000

Basic EPS=(3,050,000-88,000)/ 881000 =3.362

DILUTED EARNING PER SHARE

Numerator = net income – preferred dividends + after-tax interest savings = 3050-88+50*(1-40%) = 2992

Denominator = weighted average number of shares =shares at Jan. 1 + new shares – shares retired + exercise of options + contingent shares + conversion of bonds = (790*12/12*1.04)+(72*10/12*1.04)-(6*6/12)+14+19+20 = 934

Shares

26000

Exercise price

24

$624000

average market price

$52

Shares

12000

Exercise of options (26000-12000)

14000 shares

Contingent shares

Conversion of bonds = 500000/1000*40 = 20000

Diluted EPS=2992/934= 3.2